A gaseous hydrogen and carbon containing compound is decomposed and found to contain 85.63% C and 14.37% H by mass. The mass of 258 mLof the gas, measured at STP, was found to be 0.646 g. What is the molecular formula of the compound?
Answer #1

% by mass of Carbon = 85.63%

% by mass of Hydrogen = 14.37%

volume = 258 mL = 0.258L

at STP

P= 1 atm

T= 273K

R=0.0821L-atm/mol-K

PV=nRT

n= PV./RT = 1 x 0.258/0.0821 x 273

n = 0.0115 mole

mass = 0.646 g

number of moles = mass/molar mass

molar mass = mass/number of moles = 0.646/0.0115= 56.17 gram/mole

molar mass = 56.17 gram/mole

Element              % by mass   atomic weight      relative number                     simple ratio

C                        85.63              12.0                85.63/12.0=7.13               7.13/7.13= 1.0

H                         14.37              1.0                14.37/1.0= 14.37               14.37/7.13= 2.01= 2.0

Emperical formula = CH2

Emperical formula mass = 12+2=14

Molar mass = 56.17

n= molar mass/emperical formula mass = 56.17/14.0=4.01

n= 4.0

Molecular formula = n x emperical formula

Molecular formula = 4xCH2 = C4H8

Molecular formula = C4H8.